A sandbox 3.5 ft on each side is placed on a 79 1/3rd <--(fraction)ft by 72ft lot. Find the area of the lawn?

Found 2 solutions by Cromlix, addingup:
Answer by Cromlix(4381)

Hi there,
I assume that the sandbox is being placed
on the lawn and you want to know how much lawn
is left.
Area of lawn = 79 1/3 x 72 = 5712 ft^2
Area of sandbox = (3.5)^2 = 12.25 ft^2
5712 - 12.25 = 5699.75 ft^2 of lawn left.
Hope this helps :-)

Answer by addingup(3677)

The area of the lawn which I will call x is L*W-S^2, where S^2 is the area of the square sand box. Using the numbers in the problem:
x= (79 1/3*72)-(3.5^2). Let's solve the first parenthesis:
79 1/3 Convert this mixed number to a fraction:
(79*3)+1= 237+1= 238. So, our number is 238/3
--------------------------
238/3*72= 238/3*72/1= (238*72)/(3*1)= 5712 is the total area. Now let's deduct the sandbox:
5712-3.5^2= 5712-12.25= 5699.75 is the area of the lawn